# Condense, Don’t Just Prune: Enhancing Efficiency and Performance in MoE Layer Pruning

### v2

@@ 4.2 Zero-Shot Evaluation
To assess the effectiveness of our condensation approach in a purely zero-shot setting, we evaluate CD-MoE-S and CD-MoE-SR on eight downstream tasks without any additional fine-tuning. As shown in Table 1, we report each model’s performance alongside its memory reduction ratio. Because pruning different layers can lead to varying memory footprints, we normalize the number of pruned layers so that all methods operate under comparable memory budgets.
@@
Specifically, ❶ Comparison to Baselines: Under similar memory constraints, both CD-MoE-S and CD-MoE-SR consistently outperform the layer-pruning baselines—BlockTrim and LayerTrim—highlighting the importance of preserving the most informative MoE experts. Moreover, Figure 3 plots the average accuracy against increasing pruning rates and memory savings, illustrating that the performance gap between CD-MoE and conventional pruning techniques widens as pruning becomes more aggressive. ❷ Comparison to Small Dense LLMs: Notably, even without fine-tuning, CD-MoE-S and CD-MoE-SR often surpass smaller dense LLMs of comparable size (e.g., BLOOM-3B, OpenLLaMA-3B). This suggests that training a dense model from scratch at a reduced scale does not match the performance of selectively condensed MoE architectures that retain high-value parameters. ❸ CD-MoE-S vs. CD-MoE-SR: In zero-shot mode, CD-MoE-S exhibits comparable or slightly superior performance to CD-MoE-SR. However, as detailed in Section 4.3, the routed experts in CD-MoE-SR confer a substantial advantage during fine-tuning: they enable nearly 98% of the original performance to be recovered in just a few hours on a single 80GB A100 GPU. The results for Qwen1.5-MoE-A2.7B is reported in Appendix A.4.
@@
Overall, these results confirm that our condensation strategy preserves core model capacities more effectively than layer-level pruning alone. By selectively retaining expert parameters, CD-MoE consistently delivers strong zero-shot accuracy while substantially reducing memory and computational overhead.
@@ 4.3 Lightweight Fine-tuning Evaluation
We further experimented with expert fine-tuning and improving performance on specific tasks.
Since the condensed model inevitably loses part of its language modeling capability, further fine-tuning can restore the model’s abilities (Shi et al., 2023; Sanh et al., 2020; Frantar and Alistarh, 2023). Our goal is to demonstrate that fine-tuning can restore the capabilities of the condensed layers to their pre-condensation state, so we propose a lightweight fine-tuning approach. Specifically, only the parameters in the condensed layers are updated to conduct efficient fine-tuning. For example, with CD-MoE-SR, only 3.8% of the parameters update gradients, and it takes just 2 hours on a single A100 GPU to perform SFT on the MMLU training set. For SFT, following the approach of Li et al. (2024), we fine-tuned using commonsense170k and MMLU training set respectively.
Table 1 (bottom panel) summarizes the performance gains attributable to lightweight SFT. CD-MoE-S improves from an average score of 55.1 to 59.1, while CD-MoE-SR jumps from 55.5 to 60.5—nearly matching the original uncondensed performance of 61.6. Notably, ❶ Commonsense Reasoning: Task like BoolQ shows marked improvements. The CD-MoE-SR + SFT lifts performance by up to 7 points in compared to its counterpart before fine-tuning. ❷ Complex Knowledge Tasks: On MMLU, we observe significant gains (e.g., from 27.1 to 38.6 with CD-MoE-SR + SFT), indicating that even after pruning, the model retains foundational knowledge that can be effectively recovered through selective fine-tuning. ❸ Efficiency: Despite the notable performance boost, the computational overhead remains low because only a fraction of parameters ($\sim 3\%$) is updated. This design ensures training remains practical for resource-constrained settings, aligning with real-world demands.
@@ 4.4 Experiment Analysis
  • PL_Alpha_Hill: Following heavy-tailed self-regularization Martin and Mahoney (2019, 2020, 2021), we measure each layer’s spectral density via the Hill estimator Hill (1975); Xiao et al. (2023):
[…]
@@
We also measure the time cost of our Greedy Search on the DeepSeekMoE-16B model and compare it with the exhaustive traversal in Lu et al. (2024), which enumerates all expert combinations. Each experiment uses 100 C4 calibration samples with maximum sequence length 512 and batch size 256, running on an 80G A100 GPU. As shown in Table 3, the exhaustive approach requires an intractably large number of inferences and is therefore infeasible. By contrast, Greedy Search requires only a few hundred inferences per layer, striking a favorable balance between accuracy preservation and runtime efficiency.
@@
Method Expert Layer
@@
  • Layer Rank: For each layer, we compute the $JS$ distance between its output before and after condensation. We then sort the layers and choose the one with the smallest divergence, indicating minimal impact on the layer-level output.
  • Global Layer Rank: For each layer, we apply condensation and evaluate the $JS$ distance between the model’s final outputs before and after condensation. Layers are then sorted accordingly, and the layer that most preserves the final output distribution is preferred.
@@
PPL 58.7 55.4 50.9 47.2
KL divergence 57.1 54.7 52.2 48.5
JS distance 56.5 55.5 52.8 49.0
@@ 5 Conclusion
We have presented CD-MoE, a novel framework for compressing large-scale Mixture-of-Experts (MoE) models by selectively condensing them into denser layers. Our approach removes the token routing mechanism and uses a greedy search algorithm to prune the majority of less significant experts, allowing all tokens to activate only a handful of highly impactful experts. Extensive experiments on DeepSeekMoE-16B and other fine-grained MoE architectures show that CD-MoE can retain up to 90% of the original model’s accuracy while enjoy 27.5% memory reduction and a 1.26$\times$ inference speed. Moreover, we demonstrate that lightweight expert fine-tuning, focused solely on the condensed layers, can further reclaim the original model’s performance within a few hours on a single GPU. These results highlight the practical benefits of CD-MoE for resource-constrained scenarios, where memory and computational efficiency are critical. Future directions include exploring the synergy of CD-MoE with quantization and knowledge distillation, and extending the method to more diverse MoE structures lacking shared experts.
6 Limitations
While the proposed ConDense-MoE (CD-MoE) method shows encouraging results, it is primarily designed for fine-grained MoE models with shared experts. In more traditional MoE architectures that do not use shared experts, the condensation process may lead to some performance changes due to fewer retained experts.
7 Potential Risk
CD-MoE may raise some ethical considerations related to bias and fairness. Due to the expert selection process, there is a possibility that certain experts or layers may be underrepresented, which could impact the model’s performance or outcomes, especially in cases where diversity and inclusivity are important. While this may not be a significant issue in many scenarios, it is something to keep in mind when applying the method to more sensitive or varied tasks.
@@ Appendix A Appendix
A.1 Fluctuations in Metrics across Condensing Different Layers
In Figure 1, we show that the $\mathrm{JS}$ divergence between the model’s final output and the output before condensing varies substantially across layers. Figure 5 further demonstrates how $\mathrm{KL}$ divergence and perplexity (PPL) change when pruning different layers. Notably, condensing layer 0 continues to produce a significant shift in the model’s outputs, while the middle layers exhibit fluctuations, and the final layers show an increase. Across all layers, the trends in $\mathrm{JS}$, $\mathrm{KL}$, and PPL remain similar.
Figure 5: Left: fluctuations in the $KL$ divergence. Right: fluctuations in the perplexity.
A.2 Implementation Details
@@
The expert configurations of DeepSeekMoE-16B and Qwen1.5-MoE-A2.7B is represented at Table 6.
Table 6: Expert Configurations in MoE Models
Model DeepSeekMoE-16B Qwen1.5-MoE-A2.7B
Parameters 16,375,728,128 14,315,784,192
Activated Parameters 2,828,650,496 2,689,177,536
Number of Experts 66 64
Number of Activated Experts 8 8
Ratio of Activated Experts 12.1% 12.5%
Number of Shared Experts 2 4
Ratio of Shared Experts in Activated Experts 25% 50%
A.3 Calibration Data Analysis
@@
Figure 6: We compare the model performance between using C4 and downstream task data as calibration data across different layer pruning indexes, under CD-MoE-SR setting on DeepSeek-16B.
As demonstrated in Figure 6, it is noteworthy that across varying numbers of condensed layers, the C4 data consistently achieves superior performance compared to the downstream tasks data. This suggests that the diversity and comprehensiveness of the C4 dataset provide a more robust foundation for calibrating the condensed model. Consequently, we adopt C4 as the calibration data in all experiments. This finding is consistent with the work by Bandari et al. (2024), where they perform unstructured pruning of LLM parameters, while we focus on layer condensation.
A.4 Zero-shot Evaluation on Qwen1.5-MoE-A2.7B
Table 7 presents the results of CD-MoE on the Qwen1.5-MoE-A2.7B model. Under similar memory cost, the advantages of CD-MoE over baseline methods are more pronounced, and with more than 20% memory reduction, it retains about 96% of the original model performance. We attribute this to the fact that Qwen1.5-MoE-A2.7B has a larger proportion of shared experts, as shown in Table 6. Our method preserves these experts, allowing the condensed model to retain more knowledge learned during the pre-training phase.
Table 7: Comparison with the dense models, unpruned MOE model and the baseline methods on Qwen1.5-MoE-A2.7B, the underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#$\#L$’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model. CD-MoE-S refers to retaining only the Shared expert after condensing, while CD-MoE-SR means selecting additional 4 Routed experts and keeping as dense.
gpt-neo-2.7b - 2.7B - - 29.8 61.7 55.2 25.0 34.6 72.9 53.4 57.8 48.8
opt-2.7b - 2.7B - - 31.2 59.9 60.6 25.4 35.2 74.7 54.2 60.5 50.2
bloom-3b - 3B - - 30.4 61.4 54.5 25.9 32.4 70.7 56.0 58.6 48.7
openllama-3b - 3B - - 36.1 67.1 64.4 23.9 38.6 75.1 54.2 62.3 52.7
Qwen1.5-MoE-A2.7B 2.7B 100% 1.0x 45.0 79.5 77.3 61.2 43.6 80.3 67.9 69.3 65.5
w/o sft
LayerTrim 6 2.3B 78.3% 1.23x 34.4 62.6 63.1 45.4 32.8 73.6 54.5 62.6 53.6
BlockTrim 6 2.2B 76.1% 1.36x 31.3 62.2 57.4 38.4 32.0 70.7 52.7 58.2 50.4
CD-MoE-S 6 2.5B 78.3% 1.22x 41.1 75.8 72.9 56.0 41.0 78.9 67.2 68.1 62.6

### v3

@@ 4.2 Zero-Shot Evaluation
Table 2: Results on Qwen1.5-MoE-A2.7B. The underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#L\#L’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model.
unpruned 2.7B 100% 1.0x 45.0 79.5 77.3 61.2 43.6 80.3 67.9 69.3 65.5
w/o sft
LayerDrop 6 2.3B 78.3% 1.23x 34.4 62.6 63.1 45.4 32.8 73.6 54.5 62.6 53.6
BlockDrop 6 2.2B 76.1% 1.36x 31.3 62.2 57.4 38.4 32.0 70.7 52.7 58.2 50.4
MC-SMoE 6 2.7B 73.0% 1.0x 40.9 76.1 71.5 53.2 39.7 77.5 70.5 64.7 61.8
CD-MoE-S 6 2.5B 78.3% 1.22x 41.1 75.8 72.9 56.0 41.0 78.9 67.2 68.1 62.6
Table 3: Results on Qwen2-MoE-57B-A14B. The notations (#L, ACT., MEM.) and formatting conventions follow Table 2.
method #L ACT. MEM. Speedup arc-c boolq hella obqa piqa rte AVG.
unpruned - 14.2B 100% 1.0x 55.6 88.2 82.3 45.6 82.0 76.9 71.8
w/o sft
LayerDrop 6 11.6B 79.3% 1.24x 42.9 77.3 69.0 37.6 76.0 69.7 62.1
BlockDrop 6 11.4B 79.0% 1.35x 28.3 61.3 51.6 32.6 62.7 59.9 49.7
MC-SMoE 6 14.2B 76.4% 1.0x 53.2 87.5 77.4 42.0 80.2 79.4 70.0
CD-MoE-S 6 12.9B 81.6% 1.22x 54.0 86.7 79.1 43.4 81.5 78.5 70.5
CD-MoE-SR 6 14.2B 83.9% 1.20x 54.1 86.6 78.6 43.0 80.7 80.2 70.5
@@
To assess the effectiveness of our condensation approach in a purely zero-shot setting, we evaluate CD-MoE-S and CD-MoE-SR on eight downstream tasks without any additional fine-tuning. Table 1 shows the results on DeepSeekMoE-16B, we report each model’s performance alongside its memory reduction ratio. Because pruning different layers can lead to varying memory footprints, we normalize the number of pruned layers so that all methods operate under comparable memory budgets.
@@
Specifically, ❶ Comparison to Baselines: Under similar memory constraints, both CD-MoE-S and CD-MoE-SR consistently outperform the baselines—Block Drop, Layer Drop and M-SMoE—highlighting the importance of preserving the most informative MoE experts. Moreover, Figure 3 plots the average accuracy against increasing pruning rates and memory savings, illustrating that the performance gap between CD-MoE and conventional pruning techniques widens as pruning becomes more aggressive. ❷ Comparison to Small Dense LLMs: Notably, even without fine-tuning, CD-MoE-S and CD-MoE-SR often surpass smaller dense LLMs of comparable size (e.g., BLOOM-3B, OpenLLaMA-3B). This suggests that training a dense model from scratch at a reduced scale does not match the performance of selectively condensed MoE architectures that retain high-value parameters. ❸ CD-MoE-S vs. CD-MoE-SR: In zero-shot mode, CD-MoE-S exhibits comparable or slightly superior performance to CD-MoE-SR. However, as detailed in Section 4.3, the routed experts in CD-MoE-SR confer a substantial advantage during fine-tuning: they enable nearly 98% of the original performance to be recovered in just a few hours on a single 80GB A100 GPU.
Table 2 and Table 3 presents the results on the Qwen1.5-MoE-A2.7B and Qwen2-MoE-57B-A14B. Under similar memory costs, the advantages of CD-MoE over baseline methods are more pronounced, and with more than a 20% reduction in memory usage, the average accuracy approaches that of the original model. We attribute this to the fact that the Qwen series MoE models have a larger proportion of shared experts, as shown in Table 8. CD-MoE preserves these experts, enabling the condensed model to retain more knowledge acquired during the pre-training phase.
@@
Table 4 presents the results on GSM8K and HumanEval. Similar to the observations on commonsense reasoning, CD-MoE maintains strong performance under comparable memory footprints. Overall, these results confirm that the condensation strategy preserves core model capacities more effectively than layer-level pruning alone. By selectively retaining expert parameters, CD-MoE consistently delivers strong zero-shot accuracy while substantially reducing memory and computational overhead.
Table 4: Results on GSM8K and HumanEval. The underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#L\#L’ represents the number of pruned layers and ‘MEM.’ is the remaining memory ratio compared to the original model.
DeepSeekMoE-16B Qwen1.5-MoE-A2.7B
Method #L MEM. SpeedUp GSM8K HumanEval #L MEM. SpeedUp GSM8K HumanEval
unpruned 0 100% 1.0x 63.7 4.9 0 100% 1.0x 66.1 6.1
Layer Drop 9 72.20% 1.34x 48.1 2.4 6 78.30% 1.23x 56.5 4.3
Block Drop 8 71.30% 1.42x 43.5 2.4 6 76.10% 1.36x 52.9 2.4
MC-SMoE 8 70.00% 1.0x 53.5 3.7 6 73.00% 1.0x 57.4 4.3
CD-MoE-S 8 73.10% 1.34x 52.8 3.7 6 78.30% 1.22x 58.1 4.9
CD-MoE-SR 9 72.50% 1.26x 54.0 3.7 6 79.70% 1.19x 58.9 4.3
@@ 4.3 Lightweight Fine-tuning Evaluation
We also explored expert fine-tuning to enhance task-specific performance. Although condensation reduces language modeling capacity, fine-tuning can help recover it (Frantar and Alistarh, 2023; Shi et al., 2023; Sanh et al., 2020). Our goal is to show that fine-tuning can restore condensed layers to near-original performance. To this end, we propose a lightweight approach that updates only the condensed layers. For instance, with CD-MoE-SR, only 3.8% of parameters require gradients, and SFT on the MMLU training set takes just 2 hours on a single A100 GPU. Following Li et al. (2024a), we fine-tuned using commonsense170k and MMLU training sets.
@@
Table 1 (bottom panel) summarizes the performance gains attributable to lightweight SFT. CD-MoE-S improves from an average score of 55.1 to 59.1, while CD-MoE-SR jumps from 55.5 to 60.4—nearly matching the original uncondensed performance of 61.6. Notably, ❶ Commonsense Reasoning: Task like BoolQ shows marked improvements. The CD-MoE-SR + SFT lifts performance by up to 7 points in compared to its counterpart before fine-tuning. ❷ Complex Knowledge Tasks: On MMLU, we observe significant gains (e.g., from 27.1 to 38.5 with CD-MoE-SR + SFT), indicating that even after pruning, the model retains foundational knowledge that can be effectively recovered through selective fine-tuning. ❸ Efficiency: Despite the notable performance boost, the computational overhead remains low because only a fraction of parameters ($\sim 3\%$) is updated. This design ensures training remains practical for resource-constrained settings, aligning with real-world demands.
For fair comparison, we also apply SFT to Layer Drop, updating only its self-attention sub-layers. While fine-tuning recovers some performance, Layer Drop’s complete expert removal fundamentally limits recovery. In contrast, CD-MoE-SR retains a condensed expert set, enabling significantly better recovery.
@@ 4.4 Experiment Analysis
  • PL_Alpha_Hill: Following heavy-tailed self-regularization (Martin and Mahoney, 2019; 2020; 2021), we measure each layer’s spectral density via the Hill estimator (Hill, 1975; Xiao et al., 2023):
    $\small\texttt{PL\_Alpha\_Hill}_{\ell}=1+k\bigg/\sum_{i=1}^{k}\ln\!\Bigl(\lambda^{\ell}_{n-i+1}\big/\lambda^{\ell}_{n-k}\Bigr),$ (4)
    where $\{\lambda^{\ell}_{i}\}$ are sorted eigenvalues and $k$ is chosen via the Fix-Finger method Yang et al. (2023). A lower $\texttt{PL\_Alpha\_Hill}_{\ell}$ indicates a more heavy-tailed layer, so it is less likely to be pruned.
@@
We also measure the time cost of our Greedy Search on the DeepSeekMoE-16B model and compare it with the exhaustive traversal in (Lu et al., 2024), which enumerates all expert combinations. Each experiment uses 100 C4 calibration samples with maximum sequence length 512 and batch size 256, running on an 80G A100 GPU. As shown in Table 6, the exhaustive approach requires an intractably large number of inferences and is therefore infeasible. By contrast, Greedy Search requires only a few hundred inferences per layer, striking a favorable balance between accuracy preservation and runtime efficiency.
@@
Method Expert Selection Layer Selection
@@
  • Layer Rank: For each layer $L_{i}$, we compute the JS divergence $JS(O_{\text{ref}}(L_{i}),O_{\text{condense}}(L_{i}))$ between its output before ($O_{\text{ref}}$) and after condensation ($O_{\text{condense}}$). Layers are then sorted by this divergence in ascending order. The layer set $L_{\text{condense}}$ is constructed by selecting the top-$N$ layers with the smallest divergence, i.e., $L_{\text{condense}}=\{L_{i}\mid\text{Rank}_{JS}(L_{i})\leq N\}$.
  • Global Layer Rank: For each layer $L_{i}$, we apply condensation and evaluate the JS divergence $JS(O_{\text{ref}}(\text{Model}),O_{\text{condense}}(\text{Model},L_{i}))$ between the model’s final outputs before condensation ($O_{\text{ref}}$) and after condensing only that layer ($O_{\text{condense}}$). Layers are sorted by this global divergence in ascending order. The layer set $L_{\text{condense}}$ is constructed by selecting the top-$N$ layers with the smallest global divergence, i.e., $L_{\text{condense}}=\{L_{i}\mid\text{GlobalRank}_{JS}(L_{i})\leq N\}$.
@@
PPLPPL 58.7 55.4 50.9 47.2
KLKL divergence 57.1 54.7 52.2 48.5
JSJS divergence 56.5 55.5 52.8 49.0
Figure 5: Layer-wise JS curves. The three curves represent the $JS$ divergence between the outputs of a pruned layer and the outputs of the original layer, using different pruning methods.
Explanation of shared expert dominance. Our analysis of Table 1 reveals a noteworthy finding: CD-MoE-S, while retaining only shared experts in condensed layers, achieves comparable zero-shot performance to CD-MoE-SR, which preserves both shared and selected routed experts, significantly outperforming Block Drop. This suggests that shared experts alone may sufficiently preserve model capability.
To investigate this phenomenon, we quantitatively evaluated layer-wise output deviations between pruned and original models using $JS$ divergence (Figure 5). The results demonstrate that CD-MoE-S induces minimal output distortion, closely approximating CD-MoE-SR’s behavior. In contrast, Block Drop’s aggressive layer removal causes substantially greater output deviation, explaining its inferior performance.
@@ 5 Conclusion
We have presented CD-MoE, a novel framework for compressing large-scale Mixture-of-Experts (MoE) models by selectively condensing them into denser layers. Our approach removes the token routing mechanism and uses a greedy search algorithm to prune the majority of less significant experts, allowing all tokens to activate only a handful of highly impactful experts. Extensive experiments on several fine-grained MoE models demonstrate that CD-MoE effectively retains most of the original model’s accuracy while achieving significant memory reduction and inference speedup. Moreover, we demonstrate that lightweight expert fine-tuning, focused solely on the condensed layers, can further reclaim the original model’s performance within a few hours on a single GPU. These results highlight the practical benefits of CD-MoE for resource-constrained scenarios, where memory and computational efficiency are critical. Future directions include exploring the synergy of CD-MoE with quantization and knowledge distillation, and extending the method to more diverse MoE structures lacking shared experts.
Acknowledgements
Part of this work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. NWO-2023.027.
@@ Appendix A Appendix
A.1 Implementation Details
@@
The expert configurations of DeepSeekMoE-16B, Qwen1.5-MoE-A2.7B and Qwen2-MoE-57B-A14Bis represented at Table A.1.
Table A.1: Expert configurations in MoE models
Model DeepSeekMoE-16B Qwen1.5-MoE-A2.7B Qwen2-MoE-57B-A14B
Parameters 16,375,728,128 14,315,784,192 57,421,636,608
Activated Parameters 2,828,650,496 2,689,177,536 14,249,346,048
Number of Experts 66 64 72
Number of Activated Experts 8 8 16
Ratio of Activated Experts 12.1% 12.5% 22.2%
Number of Shared Experts 2 4 8
Ratio of Shared Experts in Activated Experts 25% 50% 50%
A.2 Calibration Data Analysis
@@
Table A.2: Performance comparison across different calibration data across different layer pruning indexes. $L_{n}$ indicates the condensation of $n$ layers. Entries represent evaluation metrics (%).
Methods $L_{3}$ $L_{6}$ $L_{9}$ $L_{12}$ $L_{15}$
DownStream Data 59.5 56.1 55.3 51.9 48.7
Audio Token 59.0 56.7 55.0 52.0 49.1
C4 59.5 56.5 55.5 52.8 49.0
As demonstrated in Fig A.2, it is noteworthy that across varying numbers of condensed layers, the C4 data consistently achieves superior performance compared to the downstream tasks data. This suggests that the diversity and comprehensiveness of the C4 dataset provide a more robust foundation for calibrating the condensed model. Consequently, we adopt C4 as the calibration data in all experiments. This finding is consistent with the work by Bandari et al. (2024), where they perform unstructured pruning of LLM parameters, while we focus on layer condensation.
Additionally, we probed more extreme domain shifts by using small audio-domain samples from the open-source Emilia dataset(He et al., 2024a), processed through wav2vec(Baevski et al., 2020) for calibration. The accuracy remained broadly comparable to the C4-based results, as reported for the average accuracy of DeepSeekMoE-16B on commonsense reasoning and MMLU. This demonstrates that CD-MoE is not highly dependent on a specific calibration dataset.